### v1

@@ Abstract
Equiangular tight frames (ETFs) are optimal packings in projective space which also yield useful decompositions of signals. Paley ETFs are constructed using number theory. In this article, we present the doubly homogeneous automorphism groups of the Paley ETFs of prime order. We also prove some properties of doubly homogeneous frames, namely that they are always ETFs and that their short circuits—i.e., dependent subsets of minimum size—form a balanced incomplete block design. We additionally characterize all $k$-homogeneous ETFs for $3\leq k\leq 5$. Finally, we revisit David Larson’s AMS Memoirs Frames, Bases, and Group Representations [HL00] coauthored with Deguang Han and Wandering Vectors for Unitary Systems and Orthogonal Wavelets [DL98] coauthored with Xingde Dai with a modern eye and focus on finite-dimensional Hilbert spaces. Keywords: equiangular tight frames, ETFs, $k$-homogeneous, $k$-transitive, Paley difference sets, $t$-designs, matroid circuits
@@ 1 Introduction
Frames are generalizations of orthonormal bases that have found wide applications in signal and image processing, compressed sensing, phase retrieval, and even characterizing emergent phenomenon in neural networks [CK13, KC07, FR13, PHD20]. Frames were originally introduced in [DS52] as a tool to loosen the definition of Fourier series. Frames over finite dimensional Hilbert spaces (see, e.g., [Wal18]), in particular so-called equiangular tight frames (ETFs) are intimately related to open problems in quantum information theory (like Zauner’s conjecture [Zau99, Zau11]) and combinatorial design theory (like the Gillespie conjecture [Gil18]). In this paper, we prove algebraic and combinatorial properties of an infinite class of ETFs (Paley ETFs). We also completely characterize all ETFs which have $3$-, $4$-, or $5$-homogeneous automorphism groups. Finally, we re-examine two of Dave Larson’s fundamental manuscripts from the turn of the millennium [HL00, DL98] in the context of frames in finite dimensions.
@@
The basics of equiangular tight frames are covered in Section 1.1, followed by sections on their symmetries Section 1.2 and linear dependencies Section 1.3. In Definition 1.22 in Section 1.3, a new combinatorial concept is introduced: the bender, which consists of the short circuits of an ETF, which are the dependent subsets of minimum size. Section 2 contains results concerning Paley ETFs, with the main result being Theorem 2.8, in which the automorphism group of an infinite class of Paley ETFs is completely characterized. Section 3 concerns $k$-homogeneous ETFs, including Theorem 3.6 which explicitly lists every $3$-, $4$-, and $5$-homogenous ETF. Finally, some key results in Dave Larson’s Memoirs [HL00, DL98] are contextualized within the framework of modern research on finite frames in Section 4. This section also includes a list of some questions and conjectures.
@@ 1.1 Equiangular Tight Frames
$A$ is called the frame bound.
In finite dimensions, any spanning set is a frame, and one often asks for additional structure, like being an equiangular tight frame. We will use the notation $\mathbb{F}$ to mean $\mathbb{R}$ or $\mathbb{C}$ throughout the paper.
@@
Let $\Phi=(\varphi_{j})_{j=1}^{n}\in\mathbb{F}^{d\times n}$. We call $\Phi$ equiangular if there exist $0\leq\alpha\leq\beta$ such that
$\left|\langle\varphi_{j},\varphi_{k}\rangle\right|=(\beta-\alpha)\delta_{j,k}+\alpha.$
@@
In $\mathbb{F}^{d}$, it is common to employ the abuse of notation using $\Phi$ to both mean the frame itself and the short fat matrix with columns the frame vectors. With that notation, $\Phi$ is a tight frame precisely when $\Phi\Phi^{\ast}=AI$. Equiangular tight frames are optimally robust to noise and up to two erasures [GKK01, SH03, HP04], are important in quantum information theory [Zau99, Zau11], and provide optimal projective codes [JKM19]. As mathematical objects which interact with a variety of fields, there are a number of conjectures about their existence or the lack thereof: Zauner’s conjecture [Zau99, Zau11], the Gillespie conjecture [Gil18], the $(d,2d)$ conjecture [FI23, IJM24], and the Fickus-Jasper conjecture [FJ19].
For many choices of $d$, $n$, $\mathbb{F}$, there does not exist an ETF. (See, e.g, [FM16].) When $n=d$, the ETFs are simply equal-norm orthogonal bases. When $n=d+1$, they span lines which go through vertices of a regular simplex.
@@
have real entries and point to vertices of a regular $d$-simplex centered at the origin. In general, one may ask if two sets of vectors are fundamentally “the same”, i.e., span the same set of lines up to a global change of basis. Fortunately, there is a relatively simple test.
@@
The test for switching equivalence has been rediscovered in different forms over many years (e.g., [Sei76, GP77, Bre90, AFF11, CW16])
@@
For $d\neq n$, there exists a $(d,n)$-ETF over $\mathbb{F}$ if and only if there exists a $(n-d,n)$-ETF over $\mathbb{F}$. Such frames are called Naimark complements of each other. Naimark complements of Parseval frames (not in general equiangular and allowing infinite dimensional Hilbert spaces) were first studied by Han and Larson [HL00], where they were called complementary frames. It was later realized that these results followed from Naimark’s dilation theorem concerning the lifting of positive operator valued measures [Neu43].
@@
Let $\Phi$ be a $(d,n)$-ETF over $\mathbb{F}$ with frame bound $A$. Then any $\Psi\in\mathbb{F}^{(n-d)\times n}$ that satisfies
@@
is an $(n-d,n)$-ETF over $\mathbb{F}$ with frame bound $A$ called a Naimark complement of $\Phi$.
Note that the inner products of distinct vectors of a Naimark complement are simply the negatives of the corresponding vectors in the first frame. Thus, by Proposition 1.5, characterizing switching equivalence of an ETF is equivalent to solving the problem for Naimark complements. Hence, it suffices to study ETFs for which $n\geq 2d$.
@@
[…]
Note that this construction via Kronecker products yields a natural way to view the rows of the Fourier matrix as being indexed lexicographically by the elements of $\mathbb{Z}_{m_{1}}\times\mathbb{Z}_{m_{2}}\times\cdots\times\mathbb{Z}_{m_{k}}$. By the fundamental theorem of finitely generated abelian groups, the character table of any finite abelian group is (up to possible permutation of entries) a Fourier matrix. Any $n\times n$ Fourier matrix is a unitary matrix scaled by $\sqrt{n}$; thus, any $d$-row all-column (i.e., $n$-column) submatrix $\Phi$ of an $n\times n$ Fourier matrix satisfies $\Phi\Phi^{\ast}=nI_{d}$. So, such submatrices always yield tight frames. We can completely characterize when such submatrices yield ETFs using combinatorial design theory (see, e.g., [CD07]).
@@
That is, the multiset of differences contains each non-identity element of $\mathbb{Z}_{7}$ exactly once. Thus $\{1,2,4\}$ is a $(7,3,1)$-difference set.
In any abelian group of size $v$, any subset of size $1$, $v-1$, or $v$ trivially forms a difference set. The difference set in Example 1.9 is the smallest non-trivial difference set. In particular, not all abelian groups have a nontrivial difference set. Examples of the construction of ETFs using difference sets appeared in [SH03, DF07, GR09], while the following theorem which completely characterizes the construction may be found in [XZG05].
@@
Let $G$ be a finite abelian group with Fourier matrix $F$. Let $\Phi$ be a all-column submatrix of $F$, with rows indexed by $X$. Then $\Phi$ is an ETF if and only if $X$ is a difference set in $G$.
@@ 1.2 The Geometric Structure of ETFs
Let $\Phi=(\varphi_{j})_{j=1}^{n}$ be a frame for $\mathbb{F}^{d}$. We call $U\in\operatorname{U}(d)$ a vector automorphism of $\Phi$ if there exists a permutation $\sigma\in S_{n}$ such that $U\varphi_{j}=\varphi_{\sigma(j)}$ for $j=1,\ldots n$. We call $U\in\operatorname{U}(d)$ a line automorphism of $\Phi$ if there exists a permutation $\sigma\in S_{n}$ such that $U\varphi_{j}\varphi_{j}^{\ast}U^{\ast}=\varphi_{\sigma(j)}\varphi_{\sigma(j)}^{\ast}$ for $j=1,\ldots n$. The group of vector automorphisms is the vector automorphism group $G$ of $\Phi$ and the quotient group of the line automorphisms modulo $\operatorname{U}(1)$ is the line automorphism group $[G]$ of $\Phi$.
By abuse of notation, we will work with elements of $\operatorname{U}(d)$ rather than $\operatorname{PU}(d)$ when analyzing the line automorphisms. In particular, $\operatorname{U}(1)$ plays no role in what permutation of the lines is induced.
@@
Let $G$ be a finite abelian group with $n\times n$ Fourier matrix $F$. Let $\Phi$ be a $d$-row, all-column submatrix of $F$, which is necessarily a tight frame. Further, define $M_{j}$ for each frame index $j$ as the $d\times d$ diagonal matrix with diagonal entries the entries of $\varphi_{j}$. Then $\{M_{j}\}_{j}$ is a group in $\operatorname{U}(d)$ which is a representation of $G$. Also, $F$ is the orbit of the all-ones vector $\mathbbm{1}$ under $\{M_{j}\}_{j}$, where $M_{j}\mathbbm{1}=\varphi_{j}$.
@@
An action of a group $G$ on a set $X=\left\{\,x_{i}\,\,\middle|\,\,i\in\{1,\ldots,n\}\,\right\}$ is $k$-transitive if for each ordered $k$-tuple with no repeated elements $(x_{i_{1}},\ldots,x_{i_{k}})$ and each permutation $\sigma\in S_{n}$, there exists a $g\in G$ such that (as ordered tuples)
@@
Clearly, $k$-transitivity implies both $(k-1)$-transitivity and $k$-homogeneity. However, $k$-homogeneity does not nest. Consider a set of size $k$ acted on by the trivial group. Then the group action is $k$-homogeneous but not $j$-homogeneous for any $1\leq j<k$.
We make a note of a result from [CW18] (cf. Proposition 1.5) that we will leverage to prove a few results and which was also the basis of the code provided by Joey Iverson that was used to test some ideas.
@@
[CW18] Let $\Phi=(\varphi_{j})_{j=1}^{n}$ be an ETF. Then there is a $g\in G$ inducing a permutation $\sigma\in S_{n}$ of the vectors if and only if for all $j\neq k$
@@
Similarly, there is a $g\in[G]$ inducing a permutation $\sigma\in S_{n}$ of the lines if and only if for all $j\neq k\neq l\neq j$
@@
Since this result relies solely on the inner products, one can see that both the vector and the line automorphism groups of an ETF and a Naimark complement of an ETF are the same.
@@ 1.3 The Combinatorial Structure of ETFs
Let $\Phi=(\varphi_{j})_{j=1}^{n}\in\mathbb{F}^{d\times n}$. The spark (or girth) of $\Phi$, $\operatorname{spark}(\Phi)$, is the size of the smallest subset of linearly dependent vectors in $\Phi$.
@@
Trivially, the largest that the spark of $n$ vectors in $\mathbb{F}^{d}$ may be is $d+1$, which happens when any subset of $d$ vectors is linear independent. When this happens, we say the vectors are full spark. Full spark frames (respectively, tight frames) of $n$ vectors in $\mathbb{F}^{d}$ form an open dense set in the space of frames (respectively, tight frames) of $n$ vectors in $\mathbb{F}^{d}$ [Cah09, ACM12, CMS17]. However, the same cannot be said when restricting to ETFs. Applying a now classical result from the compressed sensing literature [DE03] to ETFs, we have the following result.
@@
This lead to the following definition [FJKM18].
@@
If the binder is non-empty, then each of the elements of the binder (which are subsets of $\Phi$) have the same size. In certain cases (e.g., Gabor Steiner frames [BK19]), the binder forms a type of combinatorial design call a BIBD. For more on combinatorial designs see, e.g., [CD07].
@@
A $t$-$(v,k,\lambda)$-design or block design is an ordered pair $(V,\mathcal{B})$, where $V$ is a set of size $v$ and $\mathcal{B}$ is a collection of subsets of $V$ called blocks such that each block $\beta\in\mathcal{B}$ has the same size $k$ and any subset of $t$ distinct elements in $V$ is contained in precisely $\lambda$ blocks. When $t=2$, the design is called a balanced incomplete block design (BIBD).
@@
Note that $\{1,2\}$ is only in the first block, $\{1,3\}$ only in the seventh, $\{1,4\}$ only in the first, and so on. This is a $2$-$(7,3,1)$-design a.k.a. a $(7,3,1)$-BIBD.
The binder, however, of most ETFs is empty. This does not mean that the ETF is full spark, only that it does not have worst case spark. So, one may ask to study the properties of the smallest dependent sets. This leads us to introduce a new definition.
@@
Let $\Phi$ be an ETF with spark $s$. The bender111“Bender” is in honor of the smack-talking robot on Futurama, a show which is filled with math jokes. of $\Phi$ is the set of subsets of $\Phi$ (or their corresponding indices) which are size $s$ and linearly dependent. The elements of the bender are called short circuits.
In matroid theory, minimally dependent subsets of vectors, i.e., those for which any proper subset is independent, are called circuits. The size of the smallest circuit(s) is the girth, i.e., the spark. Hence the name “short circuit.”
If $\Phi$ has a non-trivial binder, then the bender is the binder and the short circuits are all simplex ETFs for their span. On the other extreme, if $\Phi$ is full spark $d+1$, then the bender is the set of all subsets of size $d+1$ of the vectors.
It was noted in [FL23] that the binder of an ETF with a doubly transitive line automorphism group is a BIBD. However, folklore in combinatorial design theory (see, e.g., [CD07, Remark 4.29]) means a much more general result holds.
@@
Let $\Phi$ be an ETF with $k$-homogeneous line automorphism group. Then the bender forms a $k$-design.
@@
Clearly, acting on a short circuit with an element of the line automorphism group results in a short circuit since unitaries do not change linear dependency. For any two different sets $X_{1}$ and $X_{2}$ of $k$ frame vectors, there is a least one element of the line automorphism group mapping $X_{1}$ to $X_{2}$, meaning that the size of the set of short circuits containing $X_{1}$ is the same as the size of the set of short circuits containing $X_{2}$. Thus, the bender is a $k$-design. ∎

### v2

@@ Abstract
We consider geometric and combinatorial characterizations of equiangular tight frames (ETFs), with the former concerning homogeneity of the vector and line symmetry groups and the latter the matroid structure. We introduce the concept of the bender of a frame, which is the collection of short circuits, which in turn are the dependent subsets of frame vectors of minimum size. We also show that ETFs with $k$-homogeneous line symmetry groups have benders which are $k$-designs. Paley ETFs are a known class of ETFs constructed using number theory. We determine the line and vector symmetry groups of all Paley ETFs and show that they are $2$-homogeneous. We additionally characterize all $k$-homogeneous ETFs for $k\geq 3$. Finally, we revisit David Larson’s AMS Memoirs Frames, Bases, and Group Representations [HL00] coauthored with Deguang Han and Wandering Vectors for Unitary Systems and Orthogonal Wavelets [DL98] coauthored with Xingde Dai with a modern eye and focus on finite-dimensional Hilbert spaces. Keywords: equiangular tight frames, ETFs, $k$-homogeneous, $k$-transitive, Paley difference sets, $t$-designs, matroid circuits MSC2020: 42C15, 20B25, 05B10, 05B35
This is dedicated to David Larson, a great mentor and researcher.
@@ 1 Introduction
Frames are generalizations of orthonormal bases that have found wide applications in signal and image processing, compressed sensing, phase retrieval, and even characterizing emergent phenomenon in neural networks [CK13, KC07, FR13, PHD20]. Frames were originally introduced in [DS52] as a tool to loosen the definition of Fourier series. Frames over finite dimensional Hilbert spaces (see, e.g., [Wal18]), in particular so-called equiangular tight frames (ETFs) are intimately related to open problems in quantum information theory (like Zauner’s conjecture [Zau99, Zau11]) and combinatorial design theory (like the Gillespie conjecture [Gil18] [^1] It is possible that this conjecture is folklore. See, e.g., p. 263 of [GR01].). In this paper, we introduce the concept of the bender of a frame, which is the collection of all linearly dependent subsets of frame vectors of minimum size (called short circuits). We prove algebraic and combinatorial properties of an infinite class of ETFs (Paley ETFs). We also characterize all ETFs which have $k$-homogeneous symmetry groups for $k\geq 3$. Finally, we re-examine two of Dave Larson’s fundamental manuscripts from the turn of the millennium [HL00, DL98] in the context of frames in finite dimensions.
@@
The basics of equiangular tight frames are covered in Section 1.1, followed by sections on their symmetries Section 1.2 and linear dependencies Section 1.3. In Definition 1.22 in Section 1.3, a new combinatorial concept is introduced: the bender, which consists of the short circuits of an ETF, which are the dependent subsets of minimum size. Section 2 contains results concerning Paley ETFs, with the main results being Theorems 2.10 and 2.19, in which the symmetry groups of all Paley ETFs are given. Section 3 concerns $k$-homogeneous ETFs, including Theorem 3.5 which characterizes $k$-homogeneous ETFs for $k\geq 3$. Finally, some key results in Dave Larson’s Memoirs [HL00, DL98] are contextualized within the framework of modern research on finite frames in Section 4. This section also includes a list of some questions and conjectures.
@@ 1.1 Equiangular Tight Frames
AA is called the frame bound.
In finite dimensions, any spanning set is a frame (in general, not tight), and one often asks for additional structure, like being an equiangular tight frame. In this section only, we will use the notation $\mathbb{F}$ (without a subscript) to mean $\mathbb{R}$ or $\mathbb{C}$.
@@
Let $\Phi=(\varphi_{j})_{j=1}^{n}\in\mathbb{F}^{d\times n}$ for $d\leq n$. We call $\Phi$ equiangular if there exist $a\leq b$ with $b>0$ such that
$\left|\langle\varphi_{j},\varphi_{k}\rangle\right|=(b-a)\delta_{j,k}+a.$
@@
We will employ the common abuse of notation using $\Phi$ to both mean the frame itself and the short fat matrix with columns the frame vectors. With that notation, $\Phi$ is a tight frame precisely when $\Phi\Phi^{\ast}=AI$. Equiangular tight frames are optimally robust to noise and up to two erasures [GKK01, SH03, HP04], are important in quantum information theory [Zau99, Zau11], and provide optimal projective codes [JKM19]. As mathematical objects which interact with a variety of fields, there are a number of conjectures about their existence or the lack thereof: Zauner’s conjecture [Zau99, Zau11], the Gillespie conjecture [Gil18], the $(d,2d)$ conjecture [FI25, IJM24], and the Fickus-Jasper conjecture [FJ19].
For many choices of $d$, $n$, $\mathbb{F}$, there does not exist an ETF. (See, e.g., [FM16].) When $n=d$, the ETFs are simply equal-norm orthogonal bases. Up to equivalence (see below) when $n=d+1$, the vectors span lines which go through vertices of a regular simplex.
@@
have real entries and point to vertices of a regular $d$-simplex centered at the origin. In general, one may ask if two sequences of vectors are fundamentally “the same”, i.e., span the same set of lines up to a global change of basis. Fortunately, there is a relatively simple test.
@@
We note that some authors also allow permutation of the indices in their definition of switching equivalence. The test for switching equivalence has been rediscovered in different forms over many years (e.g., [Sei76, GP77, Bre90, AFF11, CW16]).
@@
Every $(d,n)$-ETF over $\mathbb{F}$ has an (in general, infinitely many) associated (nd,n)(n-d,n)-ETF over $\mathbb{F}$ called a Naimark complement. Naimark complements of Parseval frames (not in general equiangular and allowing infinite dimensional Hilbert spaces) were first studied by Han and Larson [HL00], where they were called complementary frames. It was later realized that these results followed from Naimark’s dilation theorem concerning the lifting of positive operator valued measures [Neu43].
@@
Let $\Phi$ be a $(d,n)$-ETF with $d<n$ over $\mathbb{F}$ with frame bound $A$. Then any $\Psi\in\mathbb{F}^{(n-d)\times n}$ that satisfies
@@
is an (nd,n)(n-d,n)-ETF over $\mathbb{F}$ with frame bound $A$ called a Naimark complement of $\Phi$.
Note that the inner products of distinct vectors of a Naimark complement are simply the negatives of the corresponding vectors in the first frame. Thus, by Proposition 1.5, characterizing switching equivalence of an ETF is equivalent to solving the problem for Naimark complements. Hence, when studying the switching equivalence of ETFs it suffices to study ETFs with $n\geq 2d$.
@@
Let $\mathbb{Z}_{m}$ denote the integers mod $m$, and let $\zeta_{m}$ be a primitive $m$th root of unity. The Fourier matrix of $\mathbb{Z}_{m}$ is the $m\times m$ matrix with $(i,j)$-entry $\zeta^{ij}$, where indexing is $0,1,\ldots,m-1$. For $m_{1},\,m_{2},,\,\ldots,\,m_{k}\in\{2,\,3,\,\ldots\,\}$, the Fourier matrix of $\mathbb{Z}_{m_{1}}\times\mathbb{Z}_{m_{2}}\times\cdots\times\mathbb{Z}_{m_{k}}$ is the Kronecker product of the Fourier matrices of the cyclic factors.
Note that this construction via Kronecker products yields a natural way to view the rows (and columns) of the Fourier matrix as being indexed lexicographically by the elements of $\mathbb{Z}_{m_{1}}\times\mathbb{Z}_{m_{2}}\times\cdots\times\mathbb{Z}_{m_{k}}$. By the fundamental theorem of finitely generated abelian groups, the character table of any finite abelian group is (up to possible permutations of the rows and columns) a Fourier matrix. Any $n\times n$ Fourier matrix is a unitary matrix scaled by $\sqrt{n}$; thus, any $d$-row all-column (i.e., $n$-column) submatrix $\Phi$ of an $n\times n$ Fourier matrix satisfies $\Phi\Phi^{\ast}=nI_{d}$. So, such submatrices always yield tight frames. We can completely characterize when such submatrices yield ETFs using combinatorial design theory (see, e.g., [CD07]).
@@
That is, the multiset of differences contains each non-identity element of $\mathbb{Z}_{7}$ exactly once. Thus, $\{1,2,4\}$ is a $(7,3,1)$-difference set.
In any abelian group of size $v$, any subset of size $1$, v1v-1, or $v$ trivially forms a difference set. The difference set in Example 1.9 is the smallest nontrivial difference set. In particular, not all abelian groups have a nontrivial difference set. Examples of the construction of ETFs using difference sets appeared in [SH03, DF07, GR09], while the following theorem which completely characterizes the construction may be found in [XZG05].
@@
Let $G$ be a finite abelian group with Fourier matrix $F$. Let $\Phi$ be an all-column submatrix of $F$, with rows indexed by $X$. Then $\Phi$ is an ETF if and only if $X$ is a difference set in $G$.
@@ 1.2 The Geometric Structure of ETFs
Let $\Phi=(\varphi_{j})_{j=1}^{n}$ be a frame for $\mathbb{F}^{d}$. We call $U\in\operatorname{U}(d)$ a vector automorphism of $\Phi$ if there exists a permutation $\sigma\in S_{n}$ such that $U\varphi_{j}=\varphi_{\sigma(j)}$ for $j=1,\ldots n$. We call $U\in\operatorname{U}(d)$ a line automorphism of $\Phi$ if there exists a permutation $\sigma\in S_{n}$ such that $U\varphi_{j}\varphi_{j}^{\ast}U^{\ast}=\varphi_{\sigma(j)}\varphi_{\sigma(j)}^{\ast}$ for $j=1,\ldots n$. The subgroup of $S_{n}$ induced from the vector automorphisms is called the vector symmetry group $\operatorname*{Sym}_{v}(\Phi)$ of $\Phi$, and the subgroup of $S_{n}$ induced from the line automorphisms is called the line symmetry group $\operatorname*{Sym}_{\ell}(\Phi)$ of $\Phi$.
@@
Let $G$ be a finite abelian group with $n\times n$ Fourier matrix $F$. Let $\Phi$ be a $d$-row, all-column submatrix of $F$, which is necessarily a tight frame. Further, define $M^{\{j\}}$ for each frame index $j$ as the $d\times d$ diagonal matrix with diagonal entries the entries of $\varphi_{j}$. Then $\{M^{\{j\}}\}_{j}$ is a subgroup of $\operatorname{U}(d)$ which is a representation of $G$. Also, $F$ is the orbit of the all-ones vector $\mathbbm{1}$ under $\{M^{\{j\}}\}_{j}$, where $M^{\{j\}}\mathbbm{1}=\varphi_{j}$. This in turn implies that $\operatorname*{Sym}_{v}(\Phi)\leq\operatorname*{Sym}_{\ell}(\Phi)$ have a subgroup isomorphic to $G$.
@@
An action of a group $G$ on a non-empty set $X=\left\{\,x_{i}\,\,\middle|\,\,i\in\{1,\ldots,n\}\,\right\}$ is $k$-transitive if for each ordered $k$-tuple with no repeated elements $(x_{i_{1}},\ldots,x_{i_{k}})$ and each permutation $\sigma\in S_{n}$, there exists a $g\in G$ such that (as ordered tuples)
@@
Clearly, $k$-transitivity implies both (k1)(k-1)-transitivity and $k$-homogeneity. However, the relationship between $k$-homogeneity and $\ell$-homogeneity for $k\neq\ell$ is subtler. Consider a set of size $k$ acted on by the trivial group. Then the group action is $k$-homogeneous but not $j$-homogeneous for any $1\leq j<k$. On the other hand, when $k$ is less than or equal to half the size of the set acted on, $k$-homogeneity implies both (k1)(k-1)-transitivity and (k1)(k-1)-homogeneity [DM96, LW65] and in many cases also $k$-transitivity [Kan72]. In Section 2, we will prove that an infinite class at ETFs have symmetry groups which are $2$-homogeneous but not $2$-transitive.
We make a note of a result from [CW18] (cf. Proposition 1.5) that we will leverage to prove a few results and which was also the basis of the code [Ive] provided by Joey Iverson that was used to test some ideas.
@@
[CW18] Let $\Phi=(\varphi_{j})_{j=1}^{n}$ be an ETF. Then $\sigma\in\operatorname*{Sym}_{v}(\Phi)$ if and only if for all $j\neq k$
@@
Similarly, $\sigma\in\operatorname*{Sym}_{\ell}(\Phi)$ if and only if for all $j\neq k\neq l\neq j$
@@
Since this result relies solely on the inner products, one can see that the vector symmetry groups (respectively, the line symmetry groups) of an ETF and a Naimark complement of an ETF are the same.
@@ 1.3 The Combinatorial Structure of ETFs
Let $\Phi=(\varphi_{j})_{j=1}^{n}\in\mathbb{F}^{d\times n}$ for $n>d$. The spark (or girth) of $\Phi$, $\operatorname{spark}(\Phi)$, is the size of the smallest subset of linearly dependent vectors in $\Phi$.
@@
Trivially, the largest that the spark of $n>d$ vectors in $\mathbb{F}^{d}$ may be is $d+1$, which happens when any subset of $d$ vectors is linearly independent. When this happens, we say the vectors are full spark. Full spark frames (respectively, full spark tight frames) of $n$ vectors in $\mathbb{F}^{d}$ form an open dense set in the space of frames (respectively, tight frames) of $n$ vectors in $\mathbb{F}^{d}$ [Cah09, ACM12, CMS17]. However, the same cannot be said when restricting to ETFs. Applying a now classical result from the compressed sensing literature [DE03] to ETFs, we have the following result.
@@
This led to the following definition [FJKM18].
@@
If the binder is non-empty, then each of the elements of the binder (which are subsets of $\Phi$) have the same size. In certain cases (e.g., Gabor-Steiner frames [BK19]), the binder forms a type of combinatorial design call a BIBD. For more on combinatorial designs see, e.g., [CD07].
@@
A $t$-$(v,k,\lambda)$-design or block design is an ordered pair $(V,\mathcal{B})$, where $V$ is a set of size $v$ and $\mathcal{B}$ is a collection of subsets of $V$ called blocks such that each block $\beta\in\mathcal{B}$ has the same size $k$ and any subset of $t$ distinct elements in $V$ is contained in precisely $\lambda>0$ blocks. When $t=2$, the design is also called a balanced incomplete block design (BIBD).
@@
Note that $\{1,2\}$ is only in the first block, $\{1,3\}$ only in the seventh, $\{1,4\}$ only in the first, and so on. This is a 22-$(7,3,1)$-design a.k.a. a $(7,3,1)$-BIBD.
In practice we find that many ETFs have an empty binder. This does not mean that the ETF is full spark, only that it does not have worst case spark. So, one may ask to study the properties of the smallest dependent sets. This leads us to introduce a new definition.
@@
Let $\Phi$ be an ETF with spark $s$. The bender222“Bender” is in honor of the smack-talking robot on Futurama, a show which is filled with math jokes. of $\Phi$ is the set of subsets of $\Phi$ (or their corresponding indices) which are size $s$ and linearly dependent. The elements of the bender are called short circuits.
In matroid theory, minimally dependent subsets of vectors, i.e., those for which any proper subset is independent, are called circuits. The size of the smallest circuit(s) is the girth, i.e., the spark. Hence the name “short circuit.” [^3] Thanks to Matt Fickus, who suggested this name.
If $\Phi$ has a nontrivial binder, then the bender is the binder and the short circuits are all simplex ETFs for their span. On the other extreme, if $\Phi$ is full spark $d+1$, then the bender is the set of all subsets of size $d+1$ of the vectors.
It was noted in [FL25] that the binder of an ETF with a $2$-transitive line automorphism group is a BIBD. However, folklore in combinatorial design theory (see, e.g., [CD07, Remark 4.29]) means a much more general result holds.
@@
Let $\Phi$ be an ETF with $k$-homogeneous line symmetry group and spark $\geq k$. Then the bender forms a $k$-design.
@@
Note that neither scaling vectors by unimodular scalars nor performing a unitary change of basis affects linear dependency of a set. So, we will consider the short circuits to consist of the sets of the lines spanned by each of the vectors, and, thus, acting on a short circuit with a line automorphism results in a short circuit. For any two different sets $X_{1}$ and $X_{2}$ of $k$ frame vectors, there is a least one element of the line symmetry group mapping $X_{1}$ to $X_{2}$, meaning that the size of the set of short circuits containing $X_{1}$ is the same as the size of the set of short circuits containing $X_{2}$. Thus, the bender is a $k$-design. ∎